OpenStudy (jpsmarinho):

The set solution of the I equation: | (X^2-3)/X | <= 2 Is:

OpenStudy (jpsmarinho):

When I do this exercise I found the set equals to [1,3], but, I think that is not the correct answer.

OpenStudy (anonymous):

(X^2-3)/X <= 2 or (X^2-3)/X >= -2

OpenStudy (jpsmarinho):

Right,it's true ^^

OpenStudy (jpsmarinho):

But,what value you obtained?

OpenStudy (anonymous):

actually it is an "and" statement

OpenStudy (jpsmarinho):

I know that I need to separate in two cases, but the result is not [1,3] ( I found this result)

OpenStudy (jpsmarinho):

What?

OpenStudy (anonymous):

\[\frac{x^2-3}{x}\leq 2\] \[\frac{x^2-3}{x}-2\leq 0\] \[\frac{x^2-2x-3}{x}\leq 0\] \[\frac{(x-3)(x+2)}{x}\leq 0\] \(x\leq-3\) or \(0<x\leq3 \) and now you have to repeat the process on the other side

OpenStudy (jpsmarinho):

my result is different: It's wrong @satellite73 ?

OpenStudy (anonymous):

yes it is is wrong

OpenStudy (anonymous):

you have 3 factors to consider, not two they are \(x+2, x, x-3\)

OpenStudy (anonymous):

x + 2 - - - (-2) + + + + + + + + + + + + x - - - - - - - - 0 + + + + + + x-3 - - - - - - - - - - - - 3 + + + + product - - - (-2) ++ +(0) - - (3) + + +

OpenStudy (jpsmarinho):

Thanks @satellite73 ! The result is : [-3,-1] union [1,3] ^^